# Determinant and Pfaffian formulas for particle annihilation

### v1

@@ 1.2. The ghost method
The companion paper [Śni26a] introduced ghost particles for coalescence: when two particles merge, one heir and one ghost emerge, keeping the entity count at $n$ and enabling determinantal formulas.
For annihilation, the same idea applies with a twist: when two particles annihilate, an ordered pair of ghosts emerges from the collision point—invisible walkers that perform independent random walks onward, without interacting with anything. The ghosts are anonymous: only the twin pairing is preserved (Sections˜2.3 and 1.6). The entity count—survivors plus ghosts—remains exactly $n$, restoring the dimensional structure needed for determinantal methods. LABEL:fig:intro-annihilation illustrates: particles $2$ and $3$ annihilate at $c$, and a ghost pair emerges, while particles $1$ and $4$ survive.
@@ 1.3.1. The annihilation formula
Consider $n$ particles starting at $x_{1}<\cdots<x_{n}$. Suppose $k$ collisions occur, producing s=n2ks=n-2k survivors at positions $y_{1}<\cdots<y_{s}$ and $k$ ghost pairs. The ghost pairs are physically indistinguishable—they do not remember which particles created them—so we label them $1,\ldots,k$ by sampling a uniform random numbering. Ghost pair $j$ then has positions $(a_{j},b_{j})$.
@@ 1.3.2. Example: four particles, one collision
where the sign $(-1)^{I+J+1}$ comes from the Laplace expansion along columns $3$ and $4$, and each $2\times 2$ determinant is a Karlin–McGregor probability for the two surviving particles. The spatial ordering is always reversed: the originally rightward particle $J$ ends at $\min(a,b)$ and the originally leftward particle $I$ at $\max(a,b)$.
@@ 1.3.4. Relation to the LGV lemma
The annihilation formula generalizes the Karlin–McGregor / Lindström–Gessel–Viennot (LGV) determinant, just as the coalescence formula does [Śni26a]. In both cases, the proof uses a sign-reversing involution that swaps path segments at the first wrong crossing—but in annihilation, “wrong” means “crossing that violates the prescribed collision pattern” rather than any crossing at all. When no collisions occur (all particles survive, no ghosts), the formula reduces to the standard LGV determinant. See [Śni26a] for the full discussion of the LGV connection, including the relationship to Stembridge’s $D$-compatibility [Ste90] and Fisher’s vicious walkers [Fis84].
@@ 1.4. Prior work
Exact distributional results for annihilating particles have been obtained through two main approaches. Tribe and Zaboronski [TZ11] proved Pfaffian point process structure for annihilating Brownian motions under the maximal entrance law: their Pfaffian structure (Theorem 2 in [TZ11]) is exact, while their $n$-point density formula (Theorem 1) is asymptotic for large $t$. Garrod, Poplavskyi, Tribe, and Zaboronski [GPTZ18] extended the Pfaffian structure to discrete time and deterministic initial conditions. The interparticle distribution function (IPDF) method, developed by Doering and ben-Avraham [DA88] for diffusion-limited coalescence and extended to full correlation hierarchies in [Avr98], also applies to annihilation via the method of intervals [AH00]. Each of these approaches relies on analytic structure specific to the process at hand.
Our approach is combinatorial, as in the companion paper [Śni26a]: ghosts maintain a fixed entity count, enabling LGV-style determinants that apply uniformly to lattice paths, birth-death chains, and Brownian motion. Instead of the Pfaffian structure arising from infinite initial data, we have determinantal formulas that are exact for finite initial configurations. In particular, the ghost formula gives exact probabilities specifying the number of annihilations, where survivors end up, and where ghosts end up—going beyond the aggregate statistics (particle density, gap distributions) that analytic methods typically provide.
@@ 1.6. Scope and limitations
  • Continuous processes. The discrete framework (Section˜2 develops it for arbitrary planar directed acyclic graphs) extends to Brownian motion, birth-death chains, and other processes satisfying the Karlin–McGregor assumptions [KM59] (strong Markov property, meeting times are stopping times). The full continuous-time treatment is developed in the companion paper [Śni26a].
@@ 1.7. Companion papers and organization
The ghost method applies equally to coalescence ($A+A\to A$). The companion paper [Śni26a] develops the coalescence formula, a ghost-free coalescence determinant (integrating out ghost positions), and the full continuous-time treatment. The two papers share the same proof architecture—sign-reversing involution via segment swap—but differ in the ghost structure: annihilation produces ghost pairs with paired formal variables, while coalescence produces single ghosts with a staircase pattern. The Pfaffian pairwise coalescence formula (Section˜5) bridges the two models: it is proved by converting pairwise coalescence to complete annihilation via parity reclassification.
Another companion paper [Śni26] develops the ghost framework into concrete gap statistics for coalescing particles: the Rayleigh distribution for the single-gap law, an explicit joint distribution for consecutive gaps (revealing negative correlation: a large gap makes adjacent gaps smaller), a generalization of Warren’s determinantal formula [War07] from Brownian motion to all skip-free processes, and a Pfaffian empty-interval formula that recovers known results of ben-Avraham [Avr98] from the ghost framework. Voter model duality bridges the ghost framework to the Pfaffian point process theory of Tribe and Zaboronski [TZ11] and Garrod, Poplavskyi, Tribe, and Zaboronski [GPTZ18].
@@ 2. Setup
The annihilation formula holds for random walks on $\mathbb{Z}$, Brownian motion on $\mathbb{R}$, and birth-death chains on arbitrary state spaces. Following the companion paper [Śni26a], we work with spacetime graphs—a combinatorial abstraction that captures two structural properties common to all these settings: planarity (paths with swapped endpoints must cross, and non-adjacent particles cannot meet without an intermediate particle involved) and weight-preserving segment swap (exchanging path segments at crossings preserves total weight). For discrete models, the spacetime graph is literal; for continuous processes, the combinatorial structure is identical but the proof uses measure-theoretic arguments [Śni26a].
@@ 2.2. Planarity
The crossing property (P1) is Stembridge’s $D$-compatibility [Ste90]: paths with swapped endpoints must meet. The consecutive collision property (P2), introduced in the companion paper [Śni26a], ensures that non-adjacent particles cannot collide without involving intermediate ones. Both properties hold for lattice paths, random walks on $\mathbb{Z}$, birth-death chains, and Brownian motion; see [Śni26a] for verification. The consecutive collision property is essential for the sign-reversing involution: it ensures that when two paths cross, the corresponding particles are adjacent in the active set (Section˜4.6).
@@ 2.3. The annihilation model
Fix $n$ source vertices $x_{1}\prec x_{2}\prec\cdots\prec x_{n}$ in $\mathcal{X}$ and a set of target vertices in $\mathcal{Y}$. Each source vertex represents an initial particle; a path from $x_{I}$ to a target vertex $y$ represents the particle’s trajectory, with weight $W(x_{I}\to y)$ equal to the path generating function on the spacetime graph $D$.
When two particles occupy the same vertex, they annihilate: both are destroyed and a twin pair of ghosts emerges from the collision point. The ghosts do not remember which particles produced them; only the twin pairing is preserved as the ghosts drift apart.
@@ 2.5. Collision diagrams
Each initial vertex has a directed path to either a collision vertex or a survivor vertex. Distinct paths meet only at collision vertices (no crossings between collisions). The collision diagram records which collisions occur and where, but not what happens after each collision; the ghost paths are specified separately as part of the performance (Definition˜2.8).
@@ 3.1. Formal variables
Concretely, the product rule (3.1) assigns $t_{j}^{+}$ when the higher-indexed particle goes to slot $(j,1)$ and $t_{j}^{-}$ when the lower-indexed one does, so extracting $t_{j}^{+}]$ selects exactly those terms where the spatial ordering is reversed.
@@ 4. Proof of the annihilation formula
This proof parallels the coalescence proof in [Śni26a]: identical four-part structure (castings, attribution, rehearsal, sign-reversing involution) and identical involution mechanism (segment swap at the first wrong crossing). The annihilation proof is simpler because annihilation destroys both particles at a collision—no heir continues—reducing the case analysis in the involution. Nevertheless, ghost pairs differ enough from single ghosts that the success criterion, attribution map, and involution case analysis all require reworking. The present proof is self-contained; familiarity with the companion paper is not required.
@@ 4.3.1. The two-particle case
At a collision between two particles: the particle that starts more to the left continues to the ghost that ultimately ends more to the right; the particle that starts more to the right continues to the ghost that ultimately ends more to the left.
LABEL:fig:simplified-annihilation illustrates the case $\varepsilon_{j}=+1$ (that is, $a_{j}\preceq b_{j}$). Two ghost paths depart the collision vertex toward positions $a_{j}$ and $b_{j}$, see LABEL:fig:simplified-annihilation-a. The swap principle glues each incoming path to an outgoing ghost path, see LABEL:fig:simplified-annihilation-b: particle $J$ (right, larger index) is glued to the leftward ghost at $a_{j}$, and particle $I$ (left, smaller index) is glued to the rightward ghost at $b_{j}$. The case $\varepsilon_{j}=-1$ (meaning $b_{j}\prec a_{j}$) is symmetric: the labels $a_{j}$ and $b_{j}$ swap sides, so particle $J$ again goes to the leftward position and particle $I$ to the rightward one. In both cases, the spatial ordering is reversed.
@@ 4.3.3. Example
In LABEL:fig:intro-annihilation, particles $2$ and 33 ($I=2<3=J$) collide at $c$. The ghost pair has $a\preceq b$, so $\varepsilon=+1$. The swap principle glues particle 33 ($J$, right) to the leftward ghost at $a$, and particle 22 (II, left) to the rightward ghost at $b$. Survivors $1$ and $4$ pass through the collision diagram to their final positions. The resulting casting has $\pi(1)=1$ (survivor slot at $y_{1}$), $\pi(2)=(1,2)$, $\pi(3)=(1,1)$, $\pi(4)=2$ (survivor slot at $y_{2}$), with glued paths:
@@ 4.4. Rehearsal
At a collision between particles $I<J$ filling ghost pair $j$, candidacy ensures the swap principle assigns the same roles as $\pi$: if $\varepsilon_{j}=+1$, then $\pi^{-1}((j,1))>\pi^{-1}((j,2))$, so the rightward particle maps to $(j,1)$ and the leftward to (j,2)(j,2)—matching attribution. The case $\varepsilon_{j}=-1$ is symmetric. Since both particles deactivate, later collisions are unaffected. After all collisions, surviving particles retain their original paths and endpoints. Thus attribution recovers $(\pi,\mathbf{P})$. ∎
@@ 4.5. Segment swap
Property (i): after swapping at $v$, the paths $P_{I}^{\prime}$ and $P_{J}^{\prime}$ still cross at $v$. Swapping again restores the original suffixes.
Property (ii): each edge appears in exactly one path before and after the swap, so the product of weights is unchanged.
Property (iii): the bijection $\pi^{\prime}=(I\;J)\circ\pi$ differs from $\pi$ by a transposition, so $\operatorname{sgn}(\pi^{\prime})=-\operatorname{sgn}(\pi)$. ∎
@@ 5. Pairwise coalescence and Pfaffians
This section demonstrates a surprising interplay between annihilation and coalescence: we prove that the total weight of pairwise coalescence is a Pfaffian of pairwise quantities, by converting the coalescence problem to a complete annihilation problem. The annihilation formula (Section˜3.3)—developed for $A+A\to\emptyset$—turns out to be the key tool for proving a result about $A+A\to A$. This connection arises from a parity reclassification that reinterprets coalescence events as annihilations.
@@ 5.1. The coalescence model
Before deriving the Pfaffian formula, we introduce the coalescence model from the companion paper [Śni26a]. We provide only the definitions necessary for the parity reclassification.
@@ 5.1.1. Coalescence dynamics
In the coalescence model ($A+A\to A$), when two particles collide one heir and one ghost emerge: the heir continues as a visible particle while the ghost becomes an invisible walker that no longer interacts with anything. This contrasts with annihilation ($A+A\to\emptyset$), where both entities become ghosts.
@@ 5.1.2. Coalescence performance
The weight of a performance is the product of all path weights. (Unlike annihilation, coalescence ghosts are individually identifiable—each emerges from a specific junction—so no random labeling factor is needed.) See [Śni26a] for the full treatment, including the coalescence formula and the coalescence determinant.
@@ 5.2. The pairwise coalescence problem
Given $n=2k$ particles at positions $x_{1}<\cdots<x_{n}$, we ask: what is the total weight of the pairwise coalescence event—the event that by time $T$, each consecutive pair $\{x_{1},x_{2}\}$, $\{x_{3},x_{4}\}$, $\ldots$, $\{x_{n-1},x_{n}\}$ has coalesced, meaning that the particles within each pair have merged (directly or through a chain of intermediate collisions) into a single heir? (Different pairs may merge separately or may share collisions.)

### v2

@@ 1.2. The ghost method
The companion paper [Śni26] introduced ghost particles for coalescence: when two particles merge, one heir and one ghost emerge, keeping the entity count at $n$ and enabling determinantal formulas.
For annihilation, the same idea applies with a twist: when two particles annihilate, an ordered pair of ghosts emerges from the collision point—invisible walkers that perform independent random walks onward, without interacting with anything. The ghosts are anonymous: only the twin pairing is preserved (Sections˜2.3 and 1.6). The entity count—survivors plus ghosts—remains exactly $n$, restoring the dimensional structure needed for determinantal methods. We call this variant the ghost pair method. LABEL:fig:intro-annihilation illustrates: particles $2$ and $3$ annihilate at $c$, and a ghost pair emerges, while particles $1$ and $4$ survive.
@@ 1.3.1. The annihilation formula
Consider $n$ particles starting at $x_{1}\leq\cdots\leq x_{n}$. Suppose $k$ collisions occur, producing s=n2ks=n-2k survivors at positions $y_{1}<\cdots<y_{s}$ and $k$ ghost pairs. The ghost pairs are physically indistinguishable—they do not remember which particles created them—so we label them $1,\ldots,k$ by sampling a uniform random numbering. Ghost pair $j$ then has positions $(a_{j},b_{j})$.
@@ 1.3.2. Example: four particles, one collision
where the sign $(-1)^{I+J+1}$ comes from the Laplace expansion along columns $3$ and $4$, and each $2\times 2$ determinant is a Karlin–McGregor probability for the two surviving particles. The higher-indexed particle $J$ ends at $\min(a,b)$ and the lower-indexed particle $I$ at $\max(a,b)$: the index ordering is reversed relative to the ghost pair’s spatial ordering.
@@ 1.3.4. Relation to the LGV lemma
The annihilation formula generalizes the Karlin–McGregor / Lindström–Gessel–Viennot (LGV) determinant, just as the coalescence formula does [Śni26]. In both cases, the proof uses a sign-reversing involution that swaps path segments at the first wrong crossing—but in annihilation, “wrong” means “crossing that violates the prescribed collision pattern” rather than any crossing at all. When no collisions occur (all particles survive, no ghosts), the formula reduces to the standard LGV determinant. See [Śni26] for the full discussion of the LGV connection, including the relationship to Stembridge’s $D$-compatibility [Ste90] and Fisher’s vicious walkers [Fis84].
@@ 1.4. Prior work
1.4.1. Analytic approaches
Tribe and Zaboronski [TZ11] proved Pfaffian point process structure for annihilating Brownian motions under the maximal entrance law: their Pfaffian structure (Theorem 2 in [TZ11]) is exact, while their $n$-point density formula (Theorem 1) is asymptotic for large $t$. Garrod, Poplavskyi, Tribe, and Zaboronski [GPTZ18] extended this to continuous-time random walks on $\mathbb{Z}$ with spatially inhomogeneous rates and all deterministic initial conditions, covering mixed coalescence-annihilation ($\theta\in[0,1]$). Tribe and Zaboronski [TZ26] further extended this to all entrance laws. These proofs use a time-homogeneous Markov generator and ODE uniqueness.
Separately, the interparticle distribution function (IPDF) method, developed by Doering and ben-Avraham [DA88] for diffusion-limited coalescence and extended to full correlation hierarchies in [Avr98], applies to annihilation via the method of intervals [AH00].
1.4.2. The present approach
Our approach is combinatorial, as in the companion paper [Śni26]: ghosts maintain a fixed entity count, enabling LGV-style determinants that apply uniformly to lattice paths, birth-death chains, and Brownian motion, including discrete lattices with arbitrary inhomogeneous transition probabilities (varying in both space and time) where no generator is available.
The ghost formula gives exact probabilities for finite initial configurations, specifying the number of annihilations, where survivors end up, and where ghosts end up. The analytic approach gives exact correlation functions for infinite systems via a compact $2\times 2$ matrix kernel, and extends to mixed coalescence-annihilation and all entrance laws—but requires a time-homogeneous generator. Neither approach subsumes the other.
@@ 1.6. Scope and limitations
  • Continuous processes. The discrete framework (Section˜2 develops it for arbitrary planar directed acyclic graphs) extends to Brownian motion, birth-death chains, and other processes satisfying the Karlin–McGregor assumptions [KM59] (strong Markov property, meeting times are stopping times). The full continuous-time treatment is developed in the companion paper [Śni26].
@@ 1.7. Companion papers and organization
The ghost method applies equally to coalescence ($A+A\to A$). The companion paper [Śni26] develops the coalescence formula, a ghost-free coalescence determinant (integrating out ghost positions), and the full continuous-time treatment. The two papers share the same proof architecture—sign-reversing involution via segment swap—but differ in the ghost structure: annihilation produces ghost pairs with paired formal variables, while coalescence produces single ghosts with a staircase pattern. The Pfaffian pairwise coalescence formula (Section˜5) bridges the two models: it is proved by converting pairwise coalescence to complete annihilation via the cancellative labeling. A further companion paper [Śni26a] builds on this Pfaffian formula to derive an empty-interval formula for the walls (basin boundaries) of any skip-free coalescing system, together with explicit cumulants and a central limit theorem for the wall count; checkerboard duality then transfers these wall-level results to surviving particles.
@@ 2. Setup
The annihilation formula holds for random walks on $\mathbb{Z}$, Brownian motion on $\mathbb{R}$, and birth-death chains on arbitrary state spaces. Following the companion paper [Śni26], we work with spacetime graphs—a combinatorial abstraction that captures two structural properties common to all these settings: planarity (paths with swapped endpoints must cross, and non-adjacent particles cannot meet without an intermediate particle involved) and weight-preserving segment swap (exchanging path segments at crossings preserves total weight). For discrete models, the spacetime graph is literal; for continuous processes, the combinatorial structure is identical but the proof uses measure-theoretic arguments [Śni26].
@@ 2.2. Planarity
The crossing property (P1) is Stembridge’s $D$-compatibility [Ste90]: paths with swapped endpoints must meet. The consecutive collision property (P2), introduced in the companion paper [Śni26], ensures that non-adjacent particles cannot collide without involving intermediate ones. Both properties hold for lattice paths, random walks on $\mathbb{Z}$, birth-death chains, and Brownian motion; see [Śni26] for verification. The consecutive collision property is essential for the sign-reversing involution: it ensures that when two paths cross, the corresponding particles are adjacent in the active set (Section˜4.6).
@@ 2.3. The annihilation model
Fix $n$ source vertices $x_{1}\preceq x_{2}\preceq\cdots\preceq x_{n}$ in $\mathcal{X}$ and a set of target vertices in $\mathcal{Y}$. Each source vertex represents an initial particle; a path from $x_{I}$ to a target vertex $y$ represents the particle’s trajectory, with weight $W(x_{I}\to y)$ equal to the path generating function on the spacetime graph $D$.
When two particles occupy the same vertex, they annihilate: both are destroyed and a twin pair of ghosts emerges from the collision point. The ghosts do not remember which particles produced them; only the twin pairing is preserved as the ghosts drift apart. In particular, particles that share a starting position annihilate instantly: the collision occurs at time zero and only ghost paths emerge.
@@ 2.5. Collision diagrams
Each initial vertex has a directed path to either a collision vertex or a survivor vertex. Distinct paths meet only at collision vertices (no crossings between collisions). The collision diagram records which collisions occur and where, but not what happens after each collision; the ghost paths are specified separately as part of the performance (Definition˜2.8). When two initial particles share a vertex, that vertex serves simultaneously as initial vertex and collision vertex; the incoming paths have length zero.
@@ 3.1. Formal variables
Concretely, the product rule (3.1) assigns $t_{j}^{+}$ when the higher-indexed particle goes to slot $(j,1)$ and $t_{j}^{-}$ when the lower-indexed one does. Extracting $t_{j}^{+}]$ selects exactly those terms where the higher-indexed actor maps to the leftward ghost position; symmetrically, extracting $t_{j}^{-}]$ selects terms where the lower-indexed actor maps to the leftward position.
@@ 4. Proof of the annihilation formula
This proof parallels the coalescence proof in [Śni26]: identical four-part structure (castings, attribution, rehearsal, sign-reversing involution) and identical involution mechanism (segment swap at the first wrong crossing). The annihilation proof is simpler because annihilation destroys both particles at a collision—no heir continues—reducing the case analysis in the involution. Nevertheless, ghost pairs differ enough from single ghosts that the success criterion, attribution map, and involution case analysis all require reworking. The present proof is self-contained; familiarity with the companion paper is not required.
@@ 4.3.1. The two-particle case
At a collision between particles $I<J$ filling ghost pair $j$: the higher-indexed particle $J$ is glued to the ghost at the leftward position; the lower-indexed particle $I$ is glued to the ghost at the rightward position.
LABEL:fig:simplified-annihilation illustrates the case $\varepsilon_{j}=+1$ (that is, $a_{j}\preceq b_{j}$). Two ghost paths depart the collision vertex toward positions $a_{j}$ and $b_{j}$, see LABEL:fig:simplified-annihilation-a. The swap principle glues each incoming path to an outgoing ghost path, see LABEL:fig:simplified-annihilation-b: particle $J$ (higher index) is glued to the leftward ghost at $a_{j}$, and particle $I$ (lower index) is glued to the rightward ghost at $b_{j}$. The case $\varepsilon_{j}=-1$ (meaning $b_{j}\prec a_{j}$) is symmetric: the labels $a_{j}$ and $b_{j}$ swap sides, so particle $J$ again goes to the leftward position and particle $I$ to the rightward one. In both cases, the index ordering is reversed relative to the ghost pair’s spatial ordering.
@@ 4.3.3. Example
In LABEL:fig:intro-annihilation, particles $2$ and 33 ($I=2<3=J$) collide at $c$. The ghost pair has $a\preceq b$, so $\varepsilon=+1$. The swap principle glues particle 33 ($J$, higher index) to the leftward ghost at $a$, and particle 22 (II, lower index) to the rightward ghost at $b$. Survivors $1$ and $4$ pass through the collision diagram to their final positions. The resulting casting has $\pi(1)=1$ (survivor slot at $y_{1}$), $\pi(2)=(1,2)$, $\pi(3)=(1,1)$, $\pi(4)=2$ (survivor slot at $y_{2}$), with glued paths:
@@ 4.4. Rehearsal
At a collision between particles $I<J$ filling ghost pair $j$, candidacy ensures the swap principle assigns the same roles as $\pi$: if $\varepsilon_{j}=+1$, then $\pi^{-1}((j,1))>\pi^{-1}((j,2))$, so the higher-indexed particle maps to $(j,1)$ and the lower-indexed to (j,2)(j,2)—matching attribution. The case $\varepsilon_{j}=-1$ is symmetric. Since both particles deactivate, later collisions are unaffected. After all collisions, surviving particles retain their original paths and endpoints. Thus attribution recovers $(\pi,\mathbf{P})$. ∎
@@ 4.5. Segment swap
(i) After swapping at $v$, the paths $P_{I}^{\prime}$ and $P_{J}^{\prime}$ still cross at $v$. Swapping again restores the original suffixes.
(ii) Each edge appears in exactly one path before and after the swap, so the product of weights is unchanged.
(iii) The bijection $\pi^{\prime}=(I\;J)\circ\pi$ differs from $\pi$ by a transposition, so $\operatorname{sgn}(\pi^{\prime})=-\operatorname{sgn}(\pi)$. ∎
@@ 5. Pairwise coalescence and Pfaffians
This section demonstrates a surprising interplay between annihilation and coalescence: we prove that the total weight of pairwise coalescence is a Pfaffian of pairwise quantities, by converting the coalescence problem to a complete annihilation problem. The annihilation formula (Section˜3.3)—developed for $A+A\to\emptyset$—turns out to be the key tool for proving a result about $A+A\to A$. This connection arises from the cancellative labeling, which reinterprets coalescence events as annihilations.
@@ 5.1. The coalescence model
Before deriving the Pfaffian formula, we introduce the coalescence model from the companion paper [Śni26]. We provide only the definitions necessary for the cancellative labeling.
@@ 5.1.1. Coalescence dynamics
In the coalescence model ($A+A\to A$), when two particles collide one heir and one ghost emerge: the heir continues as a visible particle while the ghost drifts as an invisible walker that no longer interacts with anything. This contrasts with annihilation ($A+A\to\emptyset$), where both are destroyed and a ghost pair emerges.
@@ 5.1.2. Coalescence performance
The weight of a performance is the product of all path weights. (Unlike annihilation, coalescence ghosts are individually identifiable—each emerges from a specific junction—so no random labeling factor is needed.) See [Śni26] for the full treatment, including the coalescence formula and the coalescence determinant.
@@ 5.2. The pairwise coalescence problem
Given $n=2k$ particles at positions $x_{1}\leq\cdots\leq x_{n}$, we ask: what is the total weight of the pairwise coalescence event—the event that by time $T$, each consecutive pair $\{x_{1},x_{2}\}$, $\{x_{3},x_{4}\}$, $\ldots$, $\{x_{n-1},x_{n}\}$ has coalesced, meaning that the particles within each pair have merged (directly or through a chain of intermediate collisions) into a single heir? (Different pairs may merge separately or may share collisions.)